OpenStudy (abdela25):

step by step x- and y-intercept of the following problem? -5x+3y=-24

hartnn (hartnn):

To get x-intercept put y=0 in the equation to get y-intercept put x=0 in the equation

OpenStudy (abdela25):

i need someone to help me step by step i dont understand @hartnn

hartnn (hartnn):

-5x+3y=-24 if you put y= 0 in this, what do u get ?

OpenStudy (abdela25):

-5x+3(0) = -24

OpenStudy (abdela25):

so basically -5x=-24 because 3(0) is 0 @hartnn

hartnn (hartnn):

YES

hartnn (hartnn):

just divide both sides by -5 now

OpenStudy (abdela25):

so the answer would be x=-24/-5 which is 4.8 @hartnn

hartnn (hartnn):

x = 24/5 = 4.8 yes, correct :)

OpenStudy (abdela25):

thank you very much medal was rewarded and fanned ;) @hartnn

hartnn (hartnn):

thanks for the testimony! :D y-intercept is remaining....or did u already found out ?

OpenStudy (abdela25):

i actually already found it but can you check it for me? my final answer was y=-8 @hartnn

hartnn (hartnn):

\(\huge \checkmark \\ \huge \ddot \smile \)

OpenStudy (abdela25):

thanks!! greatly appreciated... can you help me with another word problem? @hartnn

hartnn (hartnn):

sure :)

OpenStudy (abdela25):

okay here it is: "In 2005 a Caribbean nation produced 0.7 million tons of cane sugar. Annual production was projected to decrease by 0.5 million tons each year for the next 5 years. write a linear function that models this situation."

OpenStudy (abdela25):

@hartnn

hartnn (hartnn):

so, in 2005, its 0.7 in 2006, 1 year later , its 0.2 in 2007, 2 years later, its -0.3 and so on...

OpenStudy (abdela25):

yes so the linear function would be written as f(x)....

hartnn (hartnn):

i am wondering how will they produce negative quantity of sugarcane :P

OpenStudy (abdela25):

lol maybe the company stops producing when they are negative lol

hartnn (hartnn):

do u have choices for this ?

hartnn (hartnn):

because the answer really depends on what u take for x

OpenStudy (abdela25):

no..just this word problem nothing else

OpenStudy (abdela25):

they are asking for a linear function to model the situation does that mean solve for x?

hartnn (hartnn):

ok , so i will take 'x' as number of years after 2005 so we have x= 0, y = 0.7 x=1, y = 0.2 x=2 , y = -0.3 and so on a linear function is of the form y = f(x) = mx+c

hartnn (hartnn):

where, m= slope and c = y-intercept

hartnn (hartnn):

when x =0 y = 0.7 so, 0.7 = 0+c that is c= 0.7 when x= 1, y= 0.2 0.2 = m +0.7 m = -0.5 so your linear function is y= -0.5 x + 0.7